OpenStudy (anonymous):

which one is function y=x^2+1 x^2+1

OpenStudy (anonymous):

\[y=x^2+1\] and \[x^2=y+1\]

OpenStudy (anonymous):

First of All Do u Know wat is Function

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

bt stuck there

OpenStudy (anonymous):

i think the function is y=x^2+1 (i think) im not quite sure

OpenStudy (anonymous):

wth reason

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

how hv u kn